OpenStudy (anonymous):

x + 3y = 5 -x + 6y = 4 Solve the system of equations. A) x = 1, y = 2 B) x = 2, y = 1 C) x = 1, y = 1 D) x = 0, y = 2

OpenStudy (solomonzelman):

I would suggest to add the equations together. This way you can cancel out the Xs, and you will be left with just Ys.

OpenStudy (anonymous):

I still don't get it...

OpenStudy (anonymous):

Its B

OpenStudy (anonymous):

Thank you!!! @thedomination100

OpenStudy (anonymous):

Your Welcome . Medal?

OpenStudy (camerondoherty):

@bookworm112 You give a medal by clicking best response next to a person's name. Though you cannot give out direct answers without an explanation @thedomination100 So, saying this, would you like me to show you exactly how to get to the answer?

OpenStudy (anonymous):

@camerondoherty Yes please?

OpenStudy (anonymous):

You click best response on my name

OpenStudy (camerondoherty):

Ok, I'll show you c: You have two equations: x + 3y = 5 -x + 6y = 4 So first we change up the first equation for x by subtracting 3y from both sides x = 5 - 3y Now, we can plug this in for x in the second equation: -(5 - 3y) + 6y = 4 Can you solve that?

OpenStudy (camerondoherty):

Or do you want me to solve it for you?

OpenStudy (anonymous):

I need a medal

OpenStudy (camerondoherty):

Ok, so you have -(5 - 3y) + 6y = 4 multiply everything in those parenthases by -1 -5 + 3y + 6y = 4 Now you combine like terms: 3y + 6y = 9y -5 + 9y = 4 You add 5 to both sides 9y = 9 Then you simplify to get y = 1 Now, using y = 1 you plug it in to either of the equations to find x In this case, I'm going to use the first equation: x + 3y = 5 x + 3(1) = 5 x + 3 = 5 Subtract 3 from btoh sides Then you have: x = 2 So now we have x = 2 y = 1 And we make and ordered pair which comes out to (2, 1) c:

OpenStudy (camerondoherty):

Which, as your options is B Do you get it @bookworm112 ?

OpenStudy (anonymous):

Yes, Thank you!

OpenStudy (anonymous):

@camerondoherty

OpenStudy (camerondoherty):

You're Welcome! c: @bookworm112